# Reinforcement Learning with Foundation Priors: Let the Embodied Agent Efficiently Learn on Its Own

### v3

@@ Abstract
Reinforcement learning (RL) is a promising approach for solving robotic manipulation tasks. However, it is challenging to apply the RL algorithms directly in the real world. For one thing, RL is data-intensive and typically requires millions of interactions with environments, which are impractical in real scenarios. For another, it is necessary to make heavy engineering efforts to design reward functions manually. To address these issues, we leverage foundation models in this paper. We propose Reinforcement Learning with Foundation Priors (RLFP) to utilize guidance and feedback from policy, value, and success-reward foundation models. Within this framework, we introduce the Foundation-guided Actor-Critic (FAC) algorithm, which enables embodied agents to explore more efficiently with automatic reward functions. The benefits of our framework are threefold: (1) sample efficient; (2) minimal and effective reward engineering; (3) agnostic to foundation model forms and robust to noisy priors. Our method achieves remarkable performances in various manipulation tasks on both real robots and in simulation. Across 5 dexterous tasks with real robots, FAC achieves an average success rate of 86% after one hour of real-time learning. Across 8 tasks in the simulated Meta-world, FAC achieves 100% success rates in 7/8 tasks under less than 100k frames (about 1-hour training), outperforming baseline methods with manual-designed rewards in 1M frames. We believe the RLFP framework can enable future robots to explore and learn autonomously in the physical world for more tasks.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable success in various domains, including video games [1, 2, 3], simulated robotics [4], demonstrating significant potential in solving decision-making tasks. Researchers have been increasingly applying RL algorithms to embodied agents for decades, such as real robots [5, 6, 7, 8, 9]. However, current RL algorithms encounter two primary challenges: sample efficiency and heavy reward engineering, which hinder deployment in the real world. For example, researchers [1, 10] require millions of data to master Go and Atari games, and current methods require millions of frames to solve simulated robotics tasks [11, 12, 13]. Such amounts of data are unaffordable in real. The success of current RL methods depends on manually designed rewards, which is overly burdensome. Therefore, it is essential to address these two issues.
@@
Humans are capable of acquiring skills through minimal interactions with the environment by leveraging the innate abilities and abundant commonsense accumulated in daily life. They will start from some reasonable behaviors with fewer aimless explorations, make adjustments as well as corrections, and reinforce successful behaviors. We take the baby pressing bottom as an example, which is commonly used in psychology [14, 15], to illustrate this learning paradigm. As shown in Fig. 1, a baby named Alice is presented with a novel toy box with a button. The baby observes an adult pressing the button, causing the toy box to light up and make a sound. Alice knows the rough behavior to accomplish the task, e.g., pressing the button to activate the toy. And she realizes that she can make it easier by positioning the hand closer to the button. Once she sees the toy box light up and hears the sound, she will reinforce the successful tr and repeat it in future trials.
@@
Inspired by such a learning paradigm, we explore the use of various foundation models to provide learning signals and enhance learning efficiency. Then two fundamental challenges arise: (1) what is the concrete form to present prior knowledge for RL; (2) how to leverage the corresponding prior knowledge effectively for downstream tasks. Based on how humans complete the new task shown in Fig. 1, there are three kinds of prior knowledge that answer the questions: what should I do now? am I closer to the goal? did I succeed? Notably, these prior knowledge can be aligned with the core concepts well in the Markov Decision Process (MDP), which corresponds to the policy function, the value function, and the success-reward function. Fortunately, the great success of the foundation models in natural language processing [16, 17, 18, 19] and computer vision [20, 21, 22, 23] makes it possible to acquire considerable and informative prior knowledge for embodied agents.
@@
Consequently, for systematically utilizing the abundant prior knowledge to facilitate the embodied agent efficiently learning on its own, we introduce the Reinforcement Learning from Foundation Priors (RLFP) to leverage policy, value, and success-reward prior knowledge. In this framework, the learning signals of RL are derived from corresponding foundation models. The policy prior gives a warm start behavior of the agent; the value prior informs to reach better states, and the success-reward prior gives the final success feedback.
@@
To verify the efficacy of RLFP, we instantiate an actor-critic algorithm, named Foundation-guided Actor-Critic (FAC), inspired by some works about building foundation policy or value models [24, 25, 26].. We conduct experiments on real robots and in simulation. We make extensive ablations in simulation towards each component in our framework and different qualities of the foundation models. RLFP demonstrates three key benefits: (1) Sample efficient learning. Across the 5 dexterous tasks in real robots, FAC can achieve 86% success rates on average after 1 hour of learning in real. Across the 8 tasks in the simulated Meta-world, FAC can achieve 100% success rates in 7/8 tasks under less than 100k frames (about 1 hour of training). This performance surpasses baseline methods that rely on manually designed rewards over 1M frames. (2) Minimal and effective reward engineering. The reward function is derived from the value and success-reward prior knowledge, eliminating the need for human-specified dense rewards or teleoperated demonstrations. (3) Agnostic to prior foundation model forms and robust against noisy priors. Our framework can leverage various foundation models. Notably, to ensure high efficiency and performance, we utilize several well-trained foundation models or fine-tuning foundation models. Moreover, FAC demonstrates resilience even under heavy noise or quantization errors in simulations. The contributions are:
@@ 2 Related Work
Foundation Models for Policy Learning. The ability to leverage generalized knowledge from large and varied datasets has been proved in the fields of CV and NLP. In embodied AI, researchers attempt to learn universal policies based on large language models (LLMs) or vision-language models (VLMs). Some researchers train large transformers by tokenizing inputs and inferring actions by imitation learning [27, 28, 29, 30], offline Reinforcement Learning [31], or fine-tuning the pre-trained VLMs [8] for downstream tasks. Some researchers utilize the LLMs or VLMs as reasoning tools and do low-level control based on language descriptions [32, 33, 34, 35, 36, 37, 38, 39]. The above works utilize human teleoperation to collect data for policy learning. However, it is hard to scale up human teleoperation to collect large-scale data. Some works generate code policies based on LLMs or VLMs [40, 41], while UniPi [25] predicts future videos from diffusion models [42] and generates actions via an inverse dynamics model. However, they are of poor robustness empirically because of a lack of interactions with the environments by generated codes or videos.
Foundation Models for Representation Learning. Apart from learning policies directly from the foundation models, some researchers attempt to extract universal representations for downstream tasks. Some works focus on pre-trained visual representations that initialize the perception encoder or extract latent states of image inputs [43, 44, 45, 46]. Some researchers incorporate the pre-trained LLMs or VLMs for linguistic instruction encoding [47, 48, 49]. Some have investigated how to apply the LLMs or VLMs for universal reward or value representation in RL. [50, 51, 52] build language-conditioned reward foundation models to generate task reward signals, and [24] is the first to train a universal goal-conditioned value function on large-scale unlabeled videos. Some works acquire value and success signals through human demonstrations [53, 54]. More researchers attempt to generate rewards under the value difference to avoid the design of reward functions [55, 8]. However, this learning paradigm presents certain limitations. Firstly, it lacks policy-side guidance, resulting in inefficient agent exploration. Secondly, the inferred values/rewards are continuous scalars with relatively high error and variance from the neural networks, leading to instability and difficulty in optimization. In contrast, our proposed framework leverages policy, value, and success-reward prior knowledge, which covers the abundant commonsense of solving decision-making tasks. In this way, our method is orthogonal to the current data-efficient RL algorithms [56, 11, 2, 57, 58, 59].
@@ 3 Method
This paper focuses on facilitating embodied agents to efficiently learn various tasks on their own. In Sec. 3.1, we introduce the definition and formulation of the proposed framework Reinforcement Learning with Foundation Priors (RLFP). Subsequently, in Sec. 3.2, we instantiate the Foundation-guided Actor-Critic (FAC) algorithm based on the proposed framework. Finally, we introduce how to acquire the corresponding foundation models in FAC in Sec. 3.3.
@@ 3.1 Reinforcement Learning with Foundation Priors
We model the tasks for embodied agents as the Goal-Conditioned Markov Decision Processes (GCMDP) $\mathcal{G}$: $\mathcal{G}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},{\mathcal{T}},\gamma)$. $\mathcal{S}\in\mathbb{R}^{m}$ denotes the state. $\mathcal{A}$ is the action space, which is the continuous delta movement of the end effector in this work. $\mathcal{P}$ is the transition probability function. $\mathcal{T}$ is the task identifier. $\mathcal{R}$ denotes the rewards. $\gamma$ is the discounting factor, which is equal to 0.99 in the paper. To solve the MDP, vanilla RL requires heavily reward engineering for different tasks and performs poorly under limited interactions with the environment. To enable the agent to learn efficiently and automatically, we propose the Reinforcement Learning from Foundation Priors (RLFP) framework by leveraging the policy, value, and success-reward priors.
Here we demonstrate how we formulate the corresponding priors in RLFP. Back to the case of Alice in Fig. 1, the commonsense of rough behavior can be formulated as a goal-conditioned policy function, $M_{\mathcal{\pi}}(s,\mathcal{T}):\mathcal{S}\times\mathcal{T}\rightarrow\mathcal{A}$. The prior knowledge that the state closer to the button is closer to success can be formulated as a goal-conditioned value function $M_{\mathcal{V}}(s,\mathcal{T}):\mathcal{S}\times\mathcal{T}\rightarrow\mathbb{R}^{1}$. The ability to recognize the success state can be formulated as the 0-1 success-reward function $M_{\mathcal{R}}(s,\mathcal{T}):\mathcal{S}\times\mathcal{T}\rightarrow\{0,1\}$, which equals 1 only if the task succeeds. Notably, we assume the success-reward prior is relatively precise and sound, given the simplicity of binary classification in determining success. The value and policy prior knowledge can be much noisier. Conveniently, the RLFP framework is to solve an expansion of $\mathcal{G}$, termed $\mathcal{G^{{}^{\prime}}}=(\mathcal{G},\mathcal{M})$, where $\mathcal{M}$ is the foundation model set that represents various foundation prior knowledge. Here, $M_{\mathcal{\pi}},M_{\mathcal{V}},M_{\mathcal{R}}\in\mathcal{M}$.
Compared to vanilla RL, all the signals for the RLFP should come from the foundation models. The vanilla RL relies on uninformative trial and error explorations as well as carefully and manually designed reward functions to learn the desired behaviors. It is not only of poor sample efficiency but also requires lots of human reward engineering. Instead, in RLFP, the prior knowledge from the foundation model set $\mathcal{M}$ can provide guidance or feedback in the aspects of policy, value, and success-reward, which enables automatic and more effective task resolution.
@@ 3.2 Foudation-guided Actor-Critic
Guided by Success-reward Signals. Generally, we consider the given task as MDP $\mathcal{G}_{1}$ with 0-1 success-rewards, where $\mathcal{R}_{\mathcal{G}_{1}}=M_{\mathcal{R}}(s,\mathcal{T})\in\{0,1\}$. Moreover, following the conception that human can learn from their successful trials, we propose to build a success buffer to store the previous “successful” trajectories discriminated by $M_{\mathcal{R}}$ during exploration. Each time the actor $\pi_{\phi}$ is updated through the policy gradient, it will also imitate a batch of sample data from the success buffer $\mathcal{D}_{\text{succ}}$ (if there exists data). The objective is $\mathcal{L}_{\text{succ}}(\phi)=\textbf{KL}(\pi_{\phi}(s_{t}),\mathcal{N}(a_{t},\hat{\sigma}^{2})),s_{t},a_{t}\sim\mathcal{D}_{\text{succ}}$, where $\hat{\sigma}$ is the standard deviation hyper-parameter. It allows agents to reinforce the successful trajectories. Moreover, we utilize the policy and value prior knowledge to solve the task more efficiently.
@@
Guided by Policy Regularization. To encourage the embodiment to explore the environments with the guidance of policy prior knowledge, we regularize the actor $\pi_{\phi}$ by the policy prior from $M_{\pi}(s,\mathcal{T})$ during training. Without loss of generality, we assume that the policy foundation prior follows Gaussian distributions. Thus, the regularization term is $\mathcal{L}_{\text{reg}}(\phi)=\text{KL}(\pi_{\phi},\mathcal{N}(M_{\pi}(s_{t},\mathcal{T}),\hat{\sigma}^{2}))$. Such regularization is simple to implement but effective, which is widely used in other algorithms [60, 61]. The bias caused by the policy prior can be bounded theoretically in Theorem 2.
@@
Figure 2: The overview of Foundation-guided Actor-Critic. In FAC, reward signals come from foundation success rewards and foundation value shaping. In addition to policy gradient signals, the actor is learned by the prior policy regularization and the success trajectory imitation.
@@
Guided by Reward-shaping from Value Prior. The noisy foundation policy prior might mislead agents to undesired states. We propose to guide the policy by the value model $M_{\mathcal{V}}$ to avoid unnecessary exploration of unpromising states. A natural approach is initializing with $M_{\mathcal{V}}(s,\mathcal{T})$ and fine-tuning the value functions. However, we empirically find that it performs poorly due to catastrophic forgetting issues. To better employ the value prior, we propose to utilize the reward-shaping technique [62]. We introduce the potential-based shaping function $F(s,s^{\prime},\mathcal{T})=\gamma M_{\mathcal{V}}(s^{\prime},\mathcal{T})-M_{\mathcal{V}}(s,\mathcal{T})$, where $\gamma$ is the discounting factor. Intuitively, since $M_{\mathcal{V}}$ roughly estimates the state values, $F$ can measure the increase of value towards reaching the next state $s^{\prime}$ from the current state $s$. The shaping reward will be positive if $s^{\prime}$ is better than $s$. Meanwhile, it shares the same optimal solution as the vanilla 0-1 success-reward MDP $\mathcal{G}_{1}$, under the potential-based shaping function $F$. The optimality proof and more details about reward shaping are attached in App. A.1.
@@
Foundation-guided Actor-Critic. In summary, we propose to deal with a new MDP $\mathcal{G}_{2}$, where $\mathcal{R}_{\mathcal{G}_{2}}=\lambda M_{\mathcal{R}}+F,\lambda>0$. Here, the $\lambda$ emphasizes the success feedback, which equals 100 in this work. We train the agent based on DrQ-v2 [11], a variant of Actor-Critic algorithms. We name the proposed RLFP method as Foudation-guided Actor-Critic (FAC), shown in Fig. 2. FAC learns from foundation policy guidance and the automatic reward function. Thus, the embodied agent can efficiently learn on its own based on the abundant prior knowledge. The objectives of FAC are listed in Eq. (1), where $\alpha,\beta$ as the tradeoffs are both set to 1, $y$ is the n-step TD target value, and $Q_{\bar{\theta}}$ is the target network. We employ clipped double Q-learning [63] to reduce the overestimation.
@@
We aim to study how to leverage Foundation Priors in RL, so building large-scale foundation models is out of our paper’s scope. However, we think it is an exciting future research direction. To achieve remarkable performance, we suggest well-trained or pre-trained foundation models. Here, we utilize several existing works as proxy foundation models. We choose the GPT-4V as the success-reward foundation model $M_{\mathcal{R}}$ for the real tasks. Due to the poor performance of GPT-4V in sim, we use the ground-truth 0-1 success reward in sim and distill a noisy success reward function for ablations. We utilize the VIP model [24] as the value foundation model $M_{\mathcal{V}}$, which takes in the current image observation and goal image observation to predict the value. We build the prior policy $M_{\mathcal{\pi}}$ by code generation [40, 41] or under video diffusion models [25]. We will introduce the details of each in the next section, and more details are in App. A.3.1 and App. A.3.2.

### v4

@@ Abstract
Reinforcement learning (RL) is a promising approach for solving robotic manipulation tasks. However, it is challenging to apply the RL algorithms directly in the real world. For one thing, RL is data-intensive and typically requires millions of interactions with environments, which are impractical in real scenarios. For another, it is necessary to make heavy engineering efforts to design reward functions manually. To address these issues, we leverage foundation models in this paper. We propose Reinforcement Learning with Foundation Priors (RLFP) to utilize guidance and feedback from policy, value, and success-reward foundation models. Within this framework, we introduce the Foundation-guided Actor-Critic (FAC) algorithm, which enables embodied agents to explore more efficiently with automatic reward functions. The benefits of our framework are threefold: (1) sample efficient; (2) minimal and effective reward engineering; (3) agnostic to foundation model forms and robust to noisy priors. Our method achieves remarkable performances in various manipulation tasks on both real robots and in simulation. Across 5 dexterous tasks with real robots, FAC achieves an average success rate of 86% after one hour of real-time learning. Across 8 tasks in the simulated Meta-world, FAC achieves 100% success rates in 7/8 tasks under less than 100k frames (about 1-hour training), outperforming baseline methods with manual-designed rewards in 1M frames. We believe the RLFP framework can enable future robots to explore and learn autonomously in the physical world for more tasks. Visualizations and code are available at https://yewr.github.io/rlfp.
@@ 1 Introduction
Reinforcement Learning (RL) has achieved remarkable success in various domains, including video games [1, 2, 3], simulated robotics [4], and embodied agents for decades, such as real robots [5, 6, 7, 8, 9]. However, current RL algorithms encounter two primary challenges: sample efficiency and heavy reward engineering, which hinder deployment in the real world. For example, researchers [1, 10] require millions of data to master games or solve simulated robotics tasks [11, 12, 13]. Such amounts of data are unaffordable in real. And they also depend on manually designed rewards, which is overly burdensome. Therefore, it is essential to address these two issues.
@@
Humans acquire skills through minimal interactions with the environment by leveraging the innate abilities and abundant commonsense accumulated in daily life. They start from reasonable behaviors with fewer aimless explorations, make adjustments and corrections, and reinforce successful behaviors. We take the baby pressing bottom as an example to illustrate this learning paradigm, which is commonly used in psychology [14, 15]. As shown in Fig. 1, a baby named Alice is presented with a novel toy box with a button. The baby observes an adult pressing the button, causing the box to light and make sounds. Alice knows the behavior to solve the task, e.g., pressing the button to activate the toy. She realizes that she can make it easier by positioning the hand closer to the button. Once she sees the toy box light up and hears the sound, she will reinforce the successful trial and repeat it.
@@
Inspired by such a learning paradigm, we explore various foundation models to provide learning signals and enhance learning efficiency. Then two fundamental challenges arise: (1) what is the concrete form to present prior knowledge for RL; (2) how to leverage the corresponding prior knowledge effectively for downstream tasks. Based on the example shown in Fig. 1, three kinds of prior knowledge answer the question: what should I do now? am I closer to the goal? did I succeed? These prior knowledge are aligned with the core concepts well in the Markov Decision Process (MDP), namely the policy function, the value function, and the success-reward function. Fortunately, the great success of the foundation models in natural language processing [16, 17, 18, 19] and computer vision [20, 21, 22, 23] makes it possible to acquire considerable and informative prior knowledge.
@@
Consequently, for systematically utilizing the abundant prior knowledge to facilitate the embodied agent efficiently learning on its own, we introduce the Reinforcement Learning from Foundation Priors (RLFP) to leverage policy, value, and success-reward prior knowledge. The policy prior gives a warm start behavior of the agent; the value prior informs to reach better states, and the success-reward prior gives the final success feedback.
@@
To verify the efficacy of RLFP, we instantiate an actor-critic algorithm, named Foundation-guided Actor-Critic (FAC), inspired by some works about building foundation models [24, 25, 26]. We conduct experiments on real robots and in simulation, and extensive ablations are made in simulation. RLFP demonstrates three key benefits: (1) Sample efficient learning. Across the 5 tasks on real robots, FAC can achieve 86% success rates after 1 hour of real-time learning. Across the 8 tasks in the simulated Meta-world, FAC can achieve 100% success rates in 7/8 tasks under less than 100k frames (about 1 hour of training). It surpasses baseline methods that rely on manually designed rewards over 1M frames. (2) Minimal and effective reward engineering. The reward function is derived from the value and success-reward prior knowledge, eliminating the need for human-specified dense rewards or teleoperated demonstrations. (3) Agnostic to prior foundation model forms and robust against noisy priors. FAC demonstrates resilience under quantization errors in simulations. To ensure high efficiency and performance, we utilize several well-trained foundation models or fine-tuning foundation models. The contributions are:
@@ 2 Related Work
Foundation Models for Policy Learning. The ability to leverage generalized knowledge from large and varied datasets has been proved in the fields of CV and NLP. In embodied AI, researchers attempt to learn universal policies based on large language models (LLMs) or vision-language models (VLMs). Approaches include training large transformers through imitation learning [27, 28, 29, 30], offline RL [31], or fine-tuning pre-trained VLMs for downstream tasks [8]. Others use LLMs or VLMs for reasoning and control based on language descriptions [32, 33, 34, 35, 36, 37, 38, 39]. These methods rely on human teleoperation for data collection, but scaling this is difficult. Some works generate code policies [40, 41] or predict future videos for action generation [25, 42], though they struggle with robustness due to limited interaction with environments.
Foundation Models for Representation Learning. In addition to learning policies from foundation models, some researchers focus on extracting universal representations for downstream tasks. Several works leverage pre-trained visual representations to initialize perception encoders or extract latent image states [43, 44, 45, 46], while others use pre-trained LLMs or VLMs for linguistic instruction encoding [47, 48, 49]. Researchers have also explored applying LLMs/VLMs for universal reward or value representation in RL, such as language-conditioned reward models [50, 51, 52] and universal goal-conditioned value functions trained on large-scale videos [24]. Some approaches acquire value and success signals from human demonstrations [53, 54], while others generate rewards based on value differences to avoid manual reward design [55, 8]. However, these methods face limitations, including inefficient exploration due to a lack of policy-side guidance and unstable optimization from continuous reward signals with high error. Our framework addresses these issues by foundation priors, orthogonal to existing data-efficient RL algorithms [56, 11, 2, 57, 58, 59].
@@ 3 Method
This paper aims to enable embodied agents to efficiently learn various tasks autonomously. In Sec. 3.1, we formulate the proposed Reinforcement Learning with Foundation Priors (RLFP) framework. Next, in Sec. 3.2, we present the Foundation-guided Actor-Critic (FAC) algorithm based on RLFP framework. Finally, Sec. 3.3 explains how to obtain the foundation models used in FAC.
@@ 3.1 Reinforcement Learning with Foundation Priors
We model the tasks for embodied agents as the Goal-Conditioned Markov Decision Processes (GCMDP) $\mathcal{G}$: $\mathcal{G}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},{\mathcal{T}},\gamma)$. $\mathcal{S}\in\mathbb{R}^{m}$ denotes the state. $\mathcal{A}$ is the action space, which is the continuous delta movement of the end effector in this work. $\mathcal{P}$ is the transition probability function. $\mathcal{T}$ is the task identifier. $\mathcal{R}$ denotes the rewards. $\gamma$ is the discounting factor, equal to 0.99 in the work. To learn efficiently and automatically, we propose the Reinforcement Learning from Foundation Priors (RLFP) framework by leveraging the policy, value, and success-reward priors.
Here we demonstrate how we formulate the priors in RLFP. Back to the case of Alice in Fig. 1, the commonsense of behavior can be formulated as a goal-conditioned policy function, $M_{\mathcal{\pi}}(s,\mathcal{T}):\mathcal{S}\times\mathcal{T}\rightarrow\mathcal{A}$. The prior knowledge that the state closer to the button is closer to success can be formulated as the value function $M_{\mathcal{V}}(s,\mathcal{T}):\mathcal{S}\times\mathcal{T}\rightarrow\mathbb{R}^{1}$. The ability to recognize the success state can be formulated as the 0-1 success-reward function $M_{\mathcal{R}}(s,\mathcal{T}):\mathcal{S}\times\mathcal{T}\rightarrow\{0,1\}$, which equals 1 only if the task succeeds. We assume the success-reward prior is relatively precise, given the simplicity of binary classification in determining success. The value and policy prior knowledge are noisier. The RLFP framework is to solve an expansion of $\mathcal{G}$, termed $\mathcal{G^{{}^{\prime}}}=(\mathcal{G},\mathcal{M})$, where $\mathcal{M}$ is the foundation model set that represents various foundation prior knowledge. Here, $M_{\mathcal{\pi}},M_{\mathcal{V}},M_{\mathcal{R}}\in\mathcal{M}$.
Compared to vanilla RL, all the signals for the RLFP come from the foundation models. The vanilla RL relies on uninformative trial and error explorations and manually designed reward functions. It is not only of poor sample efficiency but also requires much human reward engineering. Instead, in RLFP, prior knowledge from the foundation model set $\mathcal{M}$ provides guidance or feedback on policy, value, and success-reward, enabling more automatic and effective task resolution.
@@ 3.2 Foudation-guided Actor-Critic
Guided by Success-reward Signals. We consider the task as MDP $\mathcal{G}_{1}$ with 0-1 success rewards, where $\mathcal{R}_{\mathcal{G}_{1}}=M_{\mathcal{R}}(s,\mathcal{T})\in\{0,1\}$. Inspired by how humans learn from successful trials, we propose a success buffer to store the “successful” trajectories identified by $M_{\mathcal{R}}$. Each time the actor $\pi_{\phi}$ updates via policy gradient, it also imitates samples from the success buffer $\mathcal{D}_{\text{succ}}$ (if available). The objective is $\mathcal{L}_{\text{succ}}(\phi)=\textbf{KL}(\pi_{\phi}(s_{t}),\mathcal{N}(a_{t},\hat{\sigma}^{2})),s_{t},a_{t}\sim\mathcal{D}_{\text{succ}}$, where $\hat{\sigma}$ is the standard deviation.
@@
Guided by Policy Regularization. To encourage efficient explorations, we regularize the actor $\pi_{\phi}$ by the policy prior from $M_{\pi}(s,\mathcal{T})$. Assuming the prior follows Gaussian distributions, the regularization term is $\mathcal{L}_{\text{reg}}(\phi)=\text{KL}(\pi_{\phi},\mathcal{N}(M_{\pi}(s_{t},\mathcal{T}),\hat{\sigma}^{2}))$, which is commonly used in other algorithms [60, 61]. The bias introduced by the policy prior is bounded, shown in Theorem 2.
@@
Figure 2: The overview of Foundation-guided Actor-Critic. In FAC, rewards are derived from foundation success rewards and value shaping. Besides policy gradient updates, the actor is trained using prior policy regularization and success trajectory imitation.
@@
Guided by Reward-shaping from Value Prior. Noisy policy prior can mislead agents to undesirable states, so we propose using the value model $M_{\mathcal{V}}$ to guide exploration and avoid unpromising states. While initializing and fine-tuning with $M_{\mathcal{V}}(s,\mathcal{T})$ is a natural approach, it suffers from catastrophic forgetting. Instead, we employ the reward-shaping technique [62] using the potential-based function $F(s,s^{\prime},\mathcal{T})=\gamma M_{\mathcal{V}}(s^{\prime},\mathcal{T})-M_{\mathcal{V}}(s,\mathcal{T})$, where $\gamma$ is the discount factor. Since $M_{\mathcal{V}}$ estimates state values, $F$ measures the value increase from $s$ to $s^{\prime}$. This shaping reward is positive when $s^{\prime}$ is better than $s$ and shares the same optimal solution as the 0-1 success-reward MDP $\mathcal{G}_{1}$. Proof and details are in App. A.1.
@@
Foundation-guided Actor-Critic. In summary, we deal with a new MDP $\mathcal{G}_{2}$, where $\mathcal{R}_{\mathcal{G}_{2}}=\lambda M_{\mathcal{R}}+F$, with $\lambda$ (set to 100) emphasizing success feedback. We train the agent using DrQ-v2 [11], a variant of Actor-Critic, and call the proposed method Foundation-guided Actor-Critic (FAC). As shown in Fig. 2, FAC leverages foundation policy guidance and an automatic reward function, enabling the agent to efficiently learn from abundant prior knowledge. The objectives of FAC are detailed in Eq. (1), where tradeoff parameters $\alpha$ and $\beta$ are both set to 1, $y$ is the n-step TD target, and $Q_{\bar{\theta}}$ is the target network. We use clipped double Q-learning [63] to reduce overestimation.
@@
We focus on leveraging Foundation Priors in RL, not building large-scale foundation models, though we consider it an exciting future direction. For strong performance, we recommend using well-trained or pre-trained models. In this work, we use existing models as proxy foundation models. GPT-4V serves as the success-reward model $M_{\mathcal{R}}$ for real tasks, but due to its poor performance in simulations, we use ground-truth 0-1 success rewards in sim and distill a noisy success reward function for ablations. For value prior, we use the VIP model [24], which predicts value based on current and goal image observations. The policy prior $M_{\mathcal{\pi}}$ is built using code generation [40, 41] or video diffusion models [25]. Details are in the next section and in App. A.3.1 and App. A.3.2.